OpenStudy (anonymous):

A cylindrical vase has a diameter of 6 inches. At the bottom of the vase, there are 9 marbles, each of diameter 2 inches. The vase is filled with water up to a height of 12 inches. What is the volume of water in the vase?

OpenStudy (anonymous):

Volume = Area* Heigt... area of the base of a cylinder is a circle so A=pi*r^2 then the volume of your cylinder is pi*6^2*12 ( height of water) the volume of the water will be the volume of the cylinder minus the volume of the mables...

OpenStudy (anonymous):

so it would be v = 4/3 * pi * r3 (which is the volume of the marbles ) subtracted from pi*6^2*12?

OpenStudy (anonymous):

right but remember there are 9 marbles... so you have to multiply their volume by 9

OpenStudy (anonymous):

so i got 301.44. but my answer choices are 96π in3 27π in3 65π in3 105π in3 and none of those added up to the answer i got

OpenStudy (anonymous):

sorry my mistake, u are given the diameter of the cylinder, so the volume will be V= pi*R^2*height so V= pi*3^2*12

OpenStudy (anonymous):

i forgot to divide the diameter by 2 :D

OpenStudy (anonymous):

okay so it is 339.12-4.18 which is 334.94 and the closest answer to that is 105pi...so i guess that one?

OpenStudy (anonymous):

did you use the radius to calculate the volume of the spheres??? remember D=2*R i got that the answer is 96pi

OpenStudy (anonymous):

oh i forgot that! thanks!